OpenStudy (anonymous):

How to evaluate (6^1+6^-1)/(6^1-6^-1)

OpenStudy (anonymous):

no nice gimmick here. just write \[\frac{6+\frac{1}{6}}{6-\frac{1}{6}}\]

OpenStudy (anonymous):

multiply top and bottom by 6 to clear the fractions . you will get \[\frac{37}{35}\]

OpenStudy (anonymous):

Oh! Okay... and you get 1/6 because 6^-1 = 1/6, right?

OpenStudy (anonymous):

if you want to be fancy you can say that \[\frac{x+\frac{1}{x}}{x-\frac{1}{x}}=\frac{x^2+1}{x^2-1}\]

OpenStudy (anonymous):

yes that is right \[6^{-1}=\frac{1}{6}\]

OpenStudy (anonymous):

THanks :)